OpenStudy (anonymous):

Jack earns $50 for washing cars on the weekend. He makes an average of $4.25 in tips per hour. Write the function of Joe's earnings and solve for how much he would make after 6 hours. y = 4.25x + 50; $54.25 y = 50x + 4.25; $69.50 y = 50x + 4.25; $304.25 y = 4.25x + 50; $75.50..... helpp

OpenStudy (anonymous):

hellllllppppp ???

OpenStudy (anonymous):

X always will be with the number represent the variable that increases over time. In this case, that variable is 4.25. Since she absolutely makes 50 dollars per weekend, you don't need an x there. Can you now, get the answer? Also- On another note- if you want to figure out how much she makes working X amount of hours, put the X amount of hours in for X.

OpenStudy (anonymous):

so it's the first one?

OpenStudy (anonymous):

It is actually the last one because: Y = 4.25(6) + 50 4.25 x 6 = 25.5 25.5 + 50 = 75.50 :)

OpenStudy (anonymous):

ahhh, okay thanks so much